OpenStudy (anonymous):

Solve the quadratic equation by completing the square. 3x^2 + 12x + 39 = 0 It says I need two solutions but I'm really lost.. :l

OpenStudy (uri):

Quadratic formula,Do you know?

OpenStudy (uri):

http://blogs.discovermagazine.com/loom/files/2008/07/quadratic.jpg a:3 b:12 c:39 Can you do it now? @JennyMae18

OpenStudy (jdoe0001):

$$\bf 3x^2 + 12x + 39 = 0\\ \text{group the "x" terms first}\\ (3x^2 + 12x) + 39 = 0\\ \text{common factor of them}\\ 3(x^2+ 4x) + 39 =0\\ \text{what do we need inside the parentheses}\\ \text{ to get a perfect square trinomial?}\\ \large \color{blue}{3(x^2+ 4x+\square?) + 39 =0}\\ $$

OpenStudy (anonymous):

When I do the problem I got-2 + sqrt(-13) , 2 - sqrt(-13)

OpenStudy (uri):

Let me check. :D

OpenStudy (anonymous):

Ok :)

OpenStudy (uri):

How did you do it? Can you show? @JennyMae18

OpenStudy (anonymous):

Sure. 3x^2+12x+39=0 I subtract 39 from both sides 3x^2+12x =-39 Then I divide by 3 x^2+4x = -13 4 goes into 'c' X^2+4x+4 = -13 (x+2)^2 = -13 (x+2) = sqrt(-13) x = -2 + sqrt(-13) , -2 - sqrt (-13)

OpenStudy (uri):

Maybe you're doing this with a different formula...:3 @jdoe0001 i'm confused,help.

OpenStudy (jdoe0001):

well, I understood it as completing the square to get the "x" values

OpenStudy (jdoe0001):

$$\bf 3x^2 + 12x + 39 = 0\\ \text{group the "x" terms first}\\ (3x^2 + 12x) + 39 = 0\\ \text{common factor of them}\\ 3(x^2+ 4x) + 39 =0\\ \text{what do we need inside the parentheses}\\ \text{ to get a perfect square trinomial?}\\ \large \color{blue}{3(x^2+ 4x+\square?) + 39 =0}\\ $$

OpenStudy (jdoe0001):

so, the value we need to complete the square is \(2^2 = 4\) since is inside the parentheses, the "3" multiplies it, so we're really adding 12 keep in mind, that all we're doing is borrowing from our good friend Mr Zero, 0, so if we borrow 12, we have to give back 12 +12-12 = 0 so $$\bf 3(x^2+ 4x+2^2) + 39 =0\\ \text{so above, we added } 3\times 2^2 = 12\\ \text{so we have to substract 12}\\ 3(x+2)^2-12+39 = 0 \implies 3(x+2)^2+27=0\\ 3(x+2)^2=-27\implies (x+2)^2=-\frac{27}{3} \implies \large (x+2)^2=-9 $$

OpenStudy (anonymous):

\[x ^{2}+4x+4=-13+4=-9\] \[\left( x+2 \right)^{2}=\left( 3\iota \right)^{2}\] \[x+2=\pm3\iota, x=-2\pm3\iota \] \[x=-2+3\iota \] \[x=-2-3\iota \]